# A Novel Framework: From Trace Representation to Abnormal Trace Classification in Microservice-based systems

### v1

@@ Abstract
This paper proposes a novel framework for classifying abnormal traces into precise fault categories in Microservice-based Systems (MSS). The framework comprises two main components: MultiHAttenAE and TE-MAML. MultiHAttenAE is an unsupervised approach that constructs tailored latent trace representations for MSS using normal traces to learn the system’s typical behavior. It processes semantic, numerical, and identifier attributes from spans and logs in each trace to construct vector representations. These vectors are then utilized to create a latent representation for each trace. TE-MAML is a meta-learning-based model. It is trained using a few labeled abnormal traces from different fault categories. The representations of traces are generated via optimized MultiHAttenAE for a given microservice-based system. Once trained, TE-MAML can use the learned knowledge from previous tasks to classify new, unseen abnormal traces into novel, precise fault categories. The proposed framework is evaluated on widely used benchmark MSS with open datasets. The results demonstrate the framework’s efficacy, from Trace representation to abnormal trace classification, in employing the learned knowledge from previous tasks to effectively classify unseen abnormal traces —those that reflect abnormal patterns in distributed tracing components— into precise fault categories. This adaptability is observed both within the same MS it was initially trained on and across different MSS contexts. Our work provides a novel unsupervised method to generate tailored trace presentation for MSS, pioneers the use of AIOps to classify anomalies into precise fault categories leveraging few-shot learning, and contributes to the growing research on meta-learning for NLP.
This is some text in the document. inline, color=colorAlice!30inline, color=colorAlice!30todo: inline, color=colorAlice!30Alice: This comment is made by Alice. Another piece of text. inline, color=colorBob!30inline, color=colorBob!30todo: inline, color=colorBob!30Bob: This comment is made by Bob.
@@
Microservice, anomaly, root cause analysis, tracing, NLP, meta learning
@@ I Introduction
Microservice architecture is a software design approach where software systems are developed as a collection of small, independent services [zhou2018fault]. Traces are fundamental to understanding and monitoring Microservice-based systems (MSS) [li2022enjoy]. A trace maps the path of a user request and it is composed of interconnected spans [OpenTelemetry]. Each span is an individual operation performed by a particular service. Logs record the behaviors of each service instance in a span. Log content varies based on what the developer has decided to log.
The operational workflow for handling failures within MSS typically involves anomaly detection (AD) and root cause analysis (RCA) [soldani2022anomaly, brandon2020graph]. AD identifies deviations from normal behavior, signaling potential anomalies. Upon detection of anomalies, RCA is performed to conduct a detailed examination of the anomaly. This examination aims to pinpoint the exact service(s) causing the anomaly and to understand the underlying reasons for its occurrence. Manual AD and RCA can be time-consuming and error-prone for MSS [soldani2022anomaly] not only due to the numerous services and their intricate interactions but also the complexity and multimodality of monitoring data [soldani2022anomaly].
Artificial intelligence for IT operations (AIOps) tools place particular emphasis on trace-based AD and RCA due to their importance. Several studies focus on trace-based AD [nedelkoski2019anomaly, kaldor2017canopy, chen2021trace] represent each trace as a sequence of spans and design models to detect abnormal traces. Regarding RCA, substantial efforts [zhang2022deeptralog, chen2023tracegra, zhang2022putracead, chen2022microegrcl] have been devoted to locating the service responsible for abnormal traces for MSS, e.g., using graph neural networks. However, tooling to classify abnormal traces for MSS is still in its infancy, though the demand for such tools is increasing as they provide advantages that can significantly improve RCA. This is primarily due to the following challenges:
@@
  • Imbalanced abnormal trace distribution in MSS. This imbalance arises when some fault categories are more common than others. It poses a challenge for normal machine learning methods as they may become biased towards the more frequent fault categories [padurariu2019dealing], resulting in poor performance in identifying less frequent.
@@
Few-shot learning [wang2020generalizing] has the potential to address the challenges of imbalanced abnormal trace distribution in MSS for abnormal trace classification, as it can recognize abnormal traces from both frequent and rare fault categories by learning from a minimal number of examples. Autoencoder (AE), an unsupervised algorithm, has been utilized in trace-based AD studies [liu2020unsupervised, xie2023unsupervised, chen2023tracegra] to construct the system-specific low-dimensional trace representations (also known as latent trace representations) by fusing high-dimensional, multi-modal trace-related data, addressing the challenges driven by trace-related data and the heterogeneity of MSS. However, no studies have explored the application of few-shot learning for abnormal trace classification for MSS. Also, the use of latent trace representations generated by autoencoders for abnormal trace classification has not been explored.
@@
Given a MSS, MultiHAttenAE is trained using normal traces to construct low-dimensional trace representations tailored for this MSS. MultiHAttenAE is based on AE [hinton2006reducing], and it is in the encoder-decoder structure: an encoder that learns to fuse multi-modal high-dimensional trace-related data into latent trace representations; The decoder reconstructs the original input data from constructed trace representations. The optimized encoder can be used independently to construct low-dimensional trace representations for new, unseen traces within the MSS it is trained on.
TE-MAML uses the optimized encoder from MultiHAttenAE to construct trace representations for a given MSS. TE-MAML is built upon the meta-learning algorithm MAML [finn2017model], which allows to learn a meta-leaner with an optimal set of initial model parameters that can adapt to new tasks using only a few instances and a small number of gradient updates. These new tasks can be either from the same domain or in different domains. In our case, domains are different MSS. By leveraging MAML, TE-MAML learns from meta-training tasks from an MSS and adapts the learned knowledge (i.e., optimized initial parameters) to meta-testing tasks both within the same MSS it was initially trained on and in a different MSS. Here, each of the meta-training and meta-testing tasks is a distinct abnormal trace classification task. TE-MAML uses TE as the meta-learner, which uses a self-mechanism to weigh different parts of trace representations to identify the most relevant features for doing classification.
@@
Significance. RQ1 evaluation is critical because within-system adaptability allows the framework to maintain its effectiveness in classifying abnormal traces in the changing context. Since MSS are dynamic with frequent service updates, additions, or removals, new abnormal traces from novel fault categories may appear [zhou2018fault]. A framework that can adapt to new abnormal trace classification tasks without extensive retraining would significantly reduce the costs and rework effort while increasing practical utility in practices. RQ2 evaluates our framework’s capability to transfer learned knowledge from one MSS to another MSS with different architectures, components, and behaviors. Prior studies [chen2020logtransfer, zhang2024metalog, han2021unsupervised] have investigated cross-system adaptability in software system AD tasks to enhance the generalization capability of AIOps tools across diverse software systems. Cross-system adaptability is valuable for organizations that run several MSS, as it allows them to use the same framework for abnormal trace classification across different MSS without extensive training on each MSS. Moreover, cross-system adaptability extends the applicability of academic benchmark MSS, enabling to train on the benchmark MSS and adapt the learned knowledge to classify abnormal traces in industry-specific MSS. RQ3 identifies which parts of the framework are most critical for achieving good performance in terms of both effectiveness and efficiency. This information is valuable for others to decide on using the framework in their applications, and also they can focus on optimizing the most critical components for them.
@@
  • The practical implementation of the meta-learning algorithm MAML for a new NLP classification task: few-shot abnormal trace classification for MSS. Meta-learning has been observed actively in the domain of computer vision, while its exploration within NLP remains relatively limited [holla2020learning]
@@ II Related work and existing approaches
Existing approaches for abnormal trace classification are very rare for MSS or similar cloud-based systems. The study [nedelkoski2019anomaly2] is the only related work. It uses a convolutional neural network (CNN) to classify abnormal traces from time series-based fault categories (e.g., incremental, mean shift, gradual increase, cylinder). It characterizes the trace as a sequence of spans and uses the time-series data on span attribute “URL” to do the classification. It evaluated the model using a large dataset from a real-world production cloud. The evaluation results showed that the model can accurately classify the tested anomaly patterns. Besides, prior trace-based AD studies [nedelkoski2019anomaly, kaldor2017canopy, chen2021trace] also characterized a trace as a sequence of spans and used span attributes to construct trace representations. However, these representations might not be effective for abnormal trace classification tasks, because each span only offers insight into local behaviors for a service instance. Our study complements the prior studies and makes the novelty from these aspects:
@@ III-A Trace structure
Based on OpenTelemetry [OpenTelemetry], a trace is structured as a hierarchical tree of spans. The root span is the starting point of the trace, akin to the tree’s base, from which all other spans branch out. Each span has a parent span, except for the root span. The span that is currently being executed (i.e, active span) may contain nested sub-spans, which represent smaller units of work that are part of the larger operation encompassed by the active span to which they belong. Figure 2 shows an example tree structure for a trace, where Span A is the root span, and it triggers a sequence of calls to other spans. Figure 3 shows the same spans depicting how a request flows through the execution of each span in sequence and pinpoints the time where relevant logs are generated.
Figure 2: An example trace structure (Zhang et. al. [zhang2022deeptralog])
Figure 3: Spans and logs in the timeline (modified from Zhang et. al. [zhang2022deeptralog])
@@ III-B1 Span prepossessing and vector generation
For MSS, Span IDs are designed with a hierarchical structure reflecting the relationship between the active spans and their nested sub-spans [OpenTelemetry]. Take Figure 3 as an instance, assuming that Span A is the active span, both Span B and Span C are nested in Span A. When Span A initiates, it is assigned a Span ID “a480f2.0”, while its nested spans, Span B and Span C, are assigned with the derivative Span IDs “a480f2.1” and “a480f2.2” respectively. Spans like Span D and Span E, which are not nested sub-spans of any active spans, receive distinct Span IDs, like “a343mc.0” and “a987gq.0”, to reflect their separate execution pathways. In our study, to construct a vector representation, we abstract away the shared common prefix in Span IDs and only retain hierarchical-level digits. For example, we reassign Span A, Span B, Span C, Span D, Span E with Span ID “1.0”, “1.1”, “1.2”, “2.0”, “3.0”. We normalize Span IDs within the trace context. As such, we construct a vector representation $V_{\text{span\_id}}$ for $Span$ of $Tr$.
We concatenate textual attributes “service name” and “URL” to form a singular attribute termed “service operation”. Template-based representation of textual attributes may not be ideal given the diverse range of service operations often present in MSS [fu2022investigating, dragoni2017microservices]. Thereby, we use a neural representation method from the study [le2021log]. To construct a neural representation for service operations, we undertake three steps. Step 1. Prepossessing. We convert all uppercase letters to lowercase, substitute specific variables with standard identifiers (e.g., replace instances like “Prod1234” with “ProductID”), and remove any non-alphabetic characters. Step 2. Tokenization. We use WordPiece tokenization [wu2016google] to tokenize service operations into subwords. It starts with a vocabulary of individual characters and then incrementally combines them into more extensive sub-word units. We choose WordPiece over other methods because it gives the flexibility to handle any words (including new, unfamiliar and out-of-vocabulary ones) with a relatively small vocabulary of sub-word units, as every word is decomposable into subword units. Step 3. Neural representation. We feed the subwords, into the BERT base model [bert2018] to generate word embeddings for each sub-word. We use the word embeddings generated by the last encoding layer of the model and calculate the sentence embedding of each service operation as the average of its word embeddings. This process yields a vector representation $V_{\text{operation}}$ for service operations of $Span$ of $Tr$.
@@ III-B2 Log prepossessing and vector generation
To capture the semantic contents of each log, we extract all textual attributes: severity level (such as INFO, WARN, ERROR), component (a part of the system that generated the message), log message content (written by developers reflecting the system state). We concatenate these attributes to form a singular attribute termed “log event”. We build a neural representation for log events that omits the usual step of log parsing. This approach leverages the capability of neural representations to more effectively comprehend the semantic meanings of log events. Some past works indicate that log parsing might provide limited benefits [le2021log, hashemi2021onelog]. Our method for building a neural representation of log events follows the same three-step process as the one that is used for service operations in spans in Section III-B1. This process yields a vector representation $V_{\text{log}}\in\mathbb{R}^{d_{\text{log}}}$, which comprises the sentence embeddings of log events, for $Log$ of $Tr$. Here, $d_{\text{log}}$ represents the dimensionality of the vector space for $V_{\text{log}}$.
@@ III-B3 Trace representation construction
The encoder incorporates the multi-head attention mechanism [vaswani2017attention] , as shown in Equation 3. First, the mechanism calculates the attention each sequence element in Q should put on all other elements in K and V, where Q, K and V denote the matrices for queries, keys, and values, respectively. It computes initial attention scores by taking the dot product of Q and K, regulates these scores by ${\sqrt{d_{k}}}$ for numerical stability, and then processes the scores through a softmax function, yielding the final attention distribution. This final attention distribution assigns weights to the elements in V. Second, the mechanism characterizes each attention head ($\text{head}_{i}$), which is computed by separate learned linear projections of Q, K, and V using the matrices ${W}^{Q}$, ${W}^{K}$, ${W}^{V}$ as learnable weights. This process transforms the original matrices to produce distinct queries, keys, and values for each head, enabling a unique attention mechanism per head. Third, the mechanism details the formation of the multi-head attention, termed as MultiHead. It concatenates the outputs of all the individual attention heads into a single vector and linearly transforms this single vector using the weight matrix to produce the final representation for multihead attention representation.
@@ III-C TE-MAML for few-shot abnormal trace classification
TE-MAML is built upon the meta-learning algorithm MAML, which aims to learn a meta-leaner with an optimal set of initial model parameters that can quickly adapt to new tasks using only a few instances and a small number of gradient updates [finn2017model]. TE-MAML uses TE [vaswani2017attention] as the meta-learner. The core objective of TE-MAML is to learn the initial model parameters of this meta-learner, such that when provided with the support set S of an N-way K-shot task, it can be quickly and robustly updated to perform well on the query set Q (i.e., classify the query set Q well). Figure 5 illustrates TE-MAML’s basic architecture. TE-MAML progresses through two phases: meta-training, where the meta-leaner is trained, and meta-testing, where the meta-learner is adapted to the new tasks. We explain the meta-leaner, and meta-training and meta-testing phases in detail below.

### v3

# Few-Shot Cross-System Anomaly Trace Classification for Microservice-based systems

@@ Abstract
Microservice-based systems (MSS) may experience failures in various fault categories due to their complex and dynamic nature. To effectively handle failures, AIOps tools utilize trace-based anomaly detection and root cause analysis. In this paper, we propose a novel framework for few-shot abnormal trace classification for MSS. Our framework comprises two main components: (1) Multi-Head Attention Autoencoder for constructing system-specific trace representations, which enables (2) Transformer Encoder-based Model-Agnostic Meta-Learning to perform effective and efficient few-shot learning for abnormal trace classification. The proposed framework is evaluated on two representative MSS, Trainticket and OnlineBoutique, with open datasets. The results show that our framework can adapt the learned knowledge to classify new, unseen abnormal traces of novel fault categories both within the same system it was initially trained on and even in the different MSS. Within the same MSS, our framework achieves an average accuracy of 93.26% and 85.2% across 50 meta-testing tasks for Trainticket and OnlineBoutique, respectively, when provided with 10 instances for each task. In a cross-system context, our framework gets an average accuracy of 92.19% and 84.77% for the same meta-testing tasks of the respective system, also with 10 instances provided for each task. Our work demonstrates the applicability of achieving few-shot abnormal trace classification for MSS and shows how it can enable cross-system adaptability. This opens an avenue for building more generalized AIOps tools that require less system-specific data labeling for anomaly detection and root cause analysis.
@@
Microservice, anomaly classification, root cause analysis, tracing, NLP, meta learning, cross-system anomaly detection
@@ I Introduction
Microservice architecture is a software design approach where software systems are developed as a collection of small, independent services [1]. Traces are fundamental to understanding and monitoring Microservice-based systems (MSS) [2]. A trace maps the path of a user request and it is composed of interconnected spans [3]. Each span is an individual operation performed by a particular service. Logs record the behaviors of each service instance in a span. Log content varies based on what the developer has decided to log.
The operational workflow for handling failures within MSS typically involves anomaly detection (AD) and root cause analysis (RCA) [4, 5]. AD identifies deviations from normal behavior, signaling potential anomalies. Upon detection of anomalies, RCA is performed to conduct a detailed examination of the anomaly. This examination aims to pinpoint the exact service(s) causing the anomaly and to understand the underlying reasons for its occurrence. Manual AD and RCA can be time-consuming and error-prone for MSS [4] not only due to the numerous services and their intricate interactions but also the complexity and multimodality of monitoring data [4].
Artificial intelligence for IT operations (AIOps) tools place particular emphasis on trace-based AD and RCA due to their importance. Several studies focus on trace-based AD [6, 7, 8] represent each trace as a sequence of spans and design models to detect abnormal traces. Regarding RCA, substantial efforts [9, 10, 11, 12] have been devoted to locating the service responsible for abnormal traces for MSS, e.g., using graph neural networks. However, tooling to classify abnormal traces for MSS is still in its infancy, though the demand for such tools is increasing as they provide advantages that can significantly improve RCA. This is primarily due to the following challenges:
@@
  • Imbalanced abnormal trace distribution in MSS. This imbalance arises when some fault categories are more common than others. It poses a challenge for normal machine learning methods as they may become biased towards the more frequent fault categories [13], resulting in poor performance in identifying less frequent.
@@
Few-shot learning [14] has the potential to address the challenges of imbalanced abnormal trace distribution in MSS for abnormal trace classification, as it can recognize abnormal traces from both frequent and rare fault categories by learning from a minimal number of examples. Autoencoder (AE), an unsupervised algorithm, has been utilized in trace-based AD studies [15, 16, 10] to construct the system-specific low-dimensional trace representations (also known as latent trace representations) by fusing high-dimensional, multi-modal trace-related data, addressing the challenges driven by trace-related data and the heterogeneity of MSS. However, no studies have explored the application of few-shot learning for abnormal trace classification for MSS. Also, the use of latent trace representations generated by autoencoders for abnormal trace classification has not been explored.
@@
Given a MSS, MultiHAttenAE is trained using normal traces to construct low-dimensional trace representations tailored for this MSS. MultiHAttenAE is based on AE [17], and it is in the encoder-decoder structure: an encoder that learns to fuse multi-modal high-dimensional trace-related data into latent trace representations; The decoder reconstructs the original input data from constructed trace representations. The optimized encoder can be used independently to construct low-dimensional trace representations for new, unseen traces within the MSS it is trained on.
TE-MAML uses the optimized encoder from MultiHAttenAE to construct trace representations for a given MSS. TE-MAML is built upon the meta-learning algorithm MAML [18], which allows to learn a meta-leaner with an optimal set of initial model parameters that can adapt to new tasks using only a few instances and a small number of gradient updates. These new tasks can be either from the same domain or in different domains. In our case, domains are different MSS. By leveraging MAML, TE-MAML learns from meta-training tasks from an MSS and adapts the learned knowledge (i.e., optimized initial parameters) to meta-testing tasks both within the same MSS it was initially trained on and in a different MSS. Here, each of the meta-training and meta-testing tasks is a distinct abnormal trace classification task. TE-MAML uses TE as the meta-learner, which uses a self-mechanism to weigh different parts of trace representations to identify the most relevant features for doing classification.
@@
Significance. RQ1 evaluation is critical because within-system adaptability allows the framework to maintain its effectiveness in classifying abnormal traces in the changing context. Since MSS are dynamic with frequent service updates, additions, or removals, new abnormal traces from novel fault categories may appear [1]. A framework that can adapt to new abnormal trace classification tasks without extensive retraining would significantly reduce the costs and rework effort while increasing practical utility in practices. RQ2 evaluates our framework’s capability to transfer learned knowledge from one MSS to another MSS with different architectures, components, and behaviors. Prior studies [19, 20, 21] have investigated cross-system adaptability in software system AD tasks to enhance the generalization capability of AIOps tools across diverse software systems. Cross-system adaptability is valuable for organizations that run several MSS, as it allows them to use the same framework for abnormal trace classification across different MSS without extensive training on each MSS. Moreover, cross-system adaptability extends the applicability of academic benchmark MSS, enabling to train on the benchmark MSS and adapt the learned knowledge to classify abnormal traces in industry-specific MSS. RQ3 identifies which parts of the framework are most critical for achieving good performance in terms of both effectiveness and efficiency. This information is valuable for others to decide on using the framework in their applications, and also they can focus on optimizing the most critical components for them.
@@
  • The practical implementation of the meta-learning algorithm MAML for a new NLP classification task: few-shot abnormal trace classification for MSS. Meta-learning has been observed actively in the domain of computer vision, while its exploration within NLP remains relatively limited [22]
@@ II Related work and existing approaches
Existing approaches for abnormal trace classification are very rare for MSS or similar cloud-based systems. The study [23] is the only related work. It uses a convolutional neural network (CNN) to classify abnormal traces from time series-based fault categories (e.g., incremental, mean shift, gradual increase, cylinder). It characterizes the trace as a sequence of spans and uses the time-series data on span attribute “URL” to do the classification. It evaluated the model using a large dataset from a real-world production cloud. The evaluation results showed that the model can accurately classify the tested anomaly patterns. Besides, prior trace-based AD studies [6, 7, 8] also characterized a trace as a sequence of spans and used span attributes to construct trace representations. However, these representations might not be effective for abnormal trace classification tasks, because each span only offers insight into local behaviors for a service instance. Our study complements the prior studies and makes the novelty from these aspects:
@@ III-A Trace structure
Based on OpenTelemetry [3], a trace is structured as a hierarchical tree of spans. The root span is the starting point of the trace, akin to the tree’s base, from which all other spans branch out. Each span has a parent span, except for the root span. The span that is currently being executed (i.e, active span) may contain nested sub-spans, which represent smaller units of work that are part of the larger operation encompassed by the active span to which they belong. Figure 2 shows an example tree structure for a trace, where Span A is the root span, and it triggers a sequence of calls to other spans. Figure 3 shows the same spans depicting how a request flows through the execution of each span in sequence and pinpoints the time where relevant logs are generated.
Figure 2: An example trace structure (Zhang et. al. [9])
Figure 3: Spans and logs in the timeline (modified from Zhang et. al. [9])
@@ III-B1 Span prepossessing and vector generation
For MSS, Span IDs are designed with a hierarchical structure reflecting the relationship between the active spans and their nested sub-spans [3]. Take Figure 3 as an instance, assuming that Span A is the active span, both Span B and Span C are nested in Span A. When Span A initiates, it is assigned a Span ID “a480f2.0”, while its nested spans, Span B and Span C, are assigned with the derivative Span IDs “a480f2.1” and “a480f2.2” respectively. Spans like Span D and Span E, which are not nested sub-spans of any active spans, receive distinct Span IDs, like “a343mc.0” and “a987gq.0”, to reflect their separate execution pathways. In our study, to construct a vector representation, we abstract away the shared common prefix in Span IDs and only retain hierarchical-level digits. For example, we reassign Span A, Span B, Span C, Span D, Span E with Span ID “1.0”, “1.1”, “1.2”, “2.0”, “3.0”. We normalize Span IDs within the trace context. As such, we construct a vector representation $V_{\text{span\_id}}$ for $Span$ of $Tr$.
We concatenate textual attributes “service name” and “URL” to form a singular attribute termed “service operation”. Template-based representation of textual attributes may not be ideal given the diverse range of service operations often present in MSS [24, 25]. Thereby, we use a neural representation method from the study [26]. To construct a neural representation for service operations, we undertake three steps. Step 1. Prepossessing. We convert all uppercase letters to lowercase, substitute specific variables with standard identifiers (e.g., replace instances like “Prod1234” with “ProductID”), and remove any non-alphabetic characters. Step 2. Tokenization. We use WordPiece tokenization [27] to tokenize service operations into subwords. It starts with a vocabulary of individual characters and then incrementally combines them into more extensive sub-word units. We choose WordPiece over other methods because it gives the flexibility to handle any words (including new, unfamiliar and out-of-vocabulary ones) with a relatively small vocabulary of sub-word units, as every word is decomposable into subword units. Step 3. Neural representation. We feed the subwords, into the BERT base model [28] to generate word embeddings for each sub-word. We use the word embeddings generated by the last encoding layer of the model and calculate the sentence embedding of each service operation as the average of its word embeddings. This process yields a vector representation $V_{\text{operation}}$ for service operations of $Span$ of $Tr$.
@@ III-B2 Log prepossessing and vector generation
To capture the semantic contents of each log, we extract all textual attributes: severity level (such as INFO, WARN, ERROR), component (a part of the system that generated the message), log message content (written by developers reflecting the system state). We concatenate these attributes to form a singular attribute termed “log event”. We build a neural representation for log events that omits the usual step of log parsing. This approach leverages the capability of neural representations to more effectively comprehend the semantic meanings of log events [29]. Some past works indicate that log parsing might provide limited benefits [26, 30]. Our method for building a neural representation of log events follows the same three-step process as the one that is used for service operations in spans in Section III-B1. This process yields a vector representation $V_{\text{log}}\in\mathbb{R}^{d_{\text{log}}}$, which comprises the sentence embeddings of log events, for $Log$ of $Tr$. Here, $d_{\text{log}}$ represents the dimensionality of the vector space for $V_{\text{log}}$.
@@ III-B3 Trace representation construction
The encoder incorporates the multi-head attention mechanism [31] , as shown in Equation 3. First, the mechanism calculates the attention each sequence element in Q should put on all other elements in K and V, where Q, K and V denote the matrices for queries, keys, and values, respectively. It computes initial attention scores by taking the dot product of Q and K, regulates these scores by ${\sqrt{d_{k}}}$ for numerical stability, and then processes the scores through a softmax function, yielding the final attention distribution. This final attention distribution assigns weights to the elements in V. Second, the mechanism characterizes each attention head ($\text{head}_{i}$), which is computed by separate learned linear projections of Q, K, and V using the matrices ${W}^{Q}$, ${W}^{K}$, ${W}^{V}$ as learnable weights. This process transforms the original matrices to produce distinct queries, keys, and values for each head, enabling a unique attention mechanism per head. Third, the mechanism details the formation of the multi-head attention, termed as MultiHead. It concatenates the outputs of all the individual attention heads into a single vector and linearly transforms this single vector using the weight matrix ${W}^{O}$ to produce the final representation for multihead attention representation.
@@ III-C TE-MAML for few-shot abnormal trace classification
TE-MAML is built upon the meta-learning algorithm MAML, which aims to learn a meta-leaner with an optimal set of initial model parameters that can quickly adapt to new tasks using only a few instances and a small number of gradient updates [18]. TE-MAML uses TE [31] as the meta-learner. The core objective of TE-MAML is to learn the initial model parameters of this meta-learner, such that when provided with the support set S of an N-way K-shot task, it can be quickly and robustly updated to perform well on the query set Q (i.e., classify the query set Q well). Figure 5 illustrates TE-MAML’s basic architecture. TE-MAML progresses through two phases: meta-training, where the meta-leaner is trained, and meta-testing, where the meta-learner is adapted to the new tasks. We explain the meta-leaner, and meta-training and meta-testing phases in detail below.